OpenStudy (anonymous):

Help with this problem please. Free 2 medals will be given on your answer and your answer to other question.

OpenStudy (anonymous):

here it is

geerky42 (geerky42):

Do you know the law of sines? \[\dfrac{\sin A}{a} = \dfrac{\sin C}{c}\]

geerky42 (geerky42):

Just plug in given values then solve for c.

geerky42 (geerky42):

So all good?

OpenStudy (anonymous):

yup